A solution of malonic acid, H2C3H2O4, was standardized by titration with 0.0870 M NaOH solution. If 19.87 mL mL of the NaOH solution is required to neutralize completely 12.59 mL of the malonic acid solution, what is the molarity of the malonic acid solution?

H2C3H2O4+2NaOH→Na2C3H2O4+2H2O

nothing

M
Answer #1

Balanced chemical equation is:

2 NaOH + H2C3H2O4 ---> Na2C3H2O4 + 2 H2O

Here:

M(NaOH)=0.087 M

V(NaOH)=19.87 mL

V(H2C3H2O4)=12.59 mL

According to balanced reaction:

1*number of mol of NaOH =2*number of mol of H2C3H2O4

1*M(NaOH)*V(NaOH) =2*M(H2C3H2O4)*V(H2C3H2O4)

1*0.087*19.87 = 2*M(H2C3H2O4)*12.59

M(H2C3H2O4) = 0.0687 M

Answer: 0.0687 M